# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

### v4

@@ 3 Rank-minimizing Mechanism
if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$. That is, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00. Trivially, $f$ satisfies weak ETE. However, whether $f$ satisfies ETE or not is dependent on $q_{o_{2}}$.
@@
We assume that $r_{a_{1}}=\psi^{\prime}$ and this rank-minimizing mechanism is used. First, suppose $q_{o_{2}}=1$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=3$. Then, $r_{a_{1}}=\psi^{\prime}$ is not strategically dominated by $\psi$ and any other preference orders, because
$\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi^{\prime},\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
$\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f\left(\psi,\psi^{\prime},\psi^{\prime}\right)\right)_{a_{1}\varnothing}=\frac{1}{3}.$
@@
That is, if $p_{a_{2}}=p_{a_{3}}=\psi^{\prime}$, then $a_{1}$ is worse off by revealing $\psi$ instead of $\psi^{\prime}$. Moreover, in this case, $f$ satisfies ETE and thus this is a fair rank-minimizing mechanism.
@@
Second, suppose $q_{o_{2}}\geq 2$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=2$. Then, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$. Therefore, even if a rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy. Moreover, in this case, $f$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case. However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Therefore, in this case, $f$ does not satisfy ETE.
@@ 4 Refusal option
This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost previous studies related to ours. [^3] As exceptions, Feigenbaum et al. (2020) and Afacan (2022) consider an assignment problem where, after the initial allocation is decided, if someone refuses the assigned object, it is reallocated. Moreover, Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
@@
We reconsider strategic problems of a fair rank-minimizing mechanism. For a mechanism $f$, (revealing) $p_{a}^{\prime}$ (for $a$) is said to be strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$ for some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$.
We show that, under the uniform rank-minimizing mechanism, truth-telling is strategically dominated by another strategy when the agents have the refusal option. We let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$. First, the ranks of the truly acceptable object types are unchanged. Second, the null object type $\varnothing$ is the least preferable one under $d_{a}\in\mathcal{D}_{a}$. Thus, any object type that is less preferable than $\varnothing$ in the true preference has changed to be more preferable than $\varnothing$ in $d_{a}$. We call the strategy of agent $a$ of revealing a preference in $\mathcal{D}_{a}$ an outside option demotion strategy (hereafter ODS) of agent $a$. Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$. Suppose $q_{o_{1}}=q_{o_{2}}=1$ and $r_{a_{1}}=(o_{1},\varnothing,o_{2})$. Then, the unique ODS is revealing $d_{a_{1}}=(o_{1},o_{2},\varnothing)$. First, suppose
$p_{a_{2}}=(o_{1},o_{2},\varnothing)\text{ and }p_{a_{3}}=(o_{2},o_{1},\varnothing).$
Then,
$f^{U}\left(r_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=y=\left(\begin{array}[]{ccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right).$
@@
$f^{U}\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/2&0&1/2\\ 1/2&0&1/2\\ 0&1&0\end{array}\right)$
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$. On the other hand, suppose
$p_{a_{2}}^{\prime}=(o_{1},\varnothing,o_{2})\text{ and }p_{a_{3}}^{\prime}=(o_{1},\varnothing,o_{2}).$
Then,
$f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right)=\left(\begin{array}[]{ccc}1/3&2/3&0\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right).$
Thus, in this case, $a_{1}$ seems to be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, because $a_{1}$ prefers $\varnothing$ to $o_{2}$. However, in the model introduced in this section, $a_{1}$ can refuse $o_{2}$ and in that case, $\varnothing$ is assigned. Therefore,
$g\left(f^{U}\left(d_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)=g\left(f^{U}\left(r_{a_{1}},p_{a_{2}}^{\prime},p_{a_{3}}^{\prime}\right),r\right)\text{;}$
@@
Next, let $\hat{p}_{a_{2}}=\hat{p}_{a_{3}}=(o_{1},o_{2},\varnothing)$. Then,
$f^{U}\left(r_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
$f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
In the latter case, $a_{1}$ may refuse the assignment; that is,
$g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right),$
which is wasteful, because
$\sum\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
and $\text{and }g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
for $a^{\prime}=a_{2},a_{3}$. Thus, if
@@
then $a_{2}$ and $a_{3}$ strictly prefer and $a_{1}$ weakly prefers $\tilde{x}$ to $g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)$.
In the case where $a_{1}$ is assigned to $o_{2}$, either $a_{2}$ or $a_{3}$ is assigned to $\varnothing$ but $a_{1}$ will refuse $o_{2}$. Thus, in that case, it is more efficient that the agent who is assigned to $\varnothing$ is assigned to $o_{2}$ and $a_{1}$ is assigned to $\varnothing$. That is, revealing $d_{a_{1}}$ and refusing some unacceptable assignment may make $a_{1}$ better off but result in a wasteful assignment, because some agents may desire the object type refused by $a_{1}$.
Note that $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ can be a Nash equilibrium when $r_{a_{2}}=\hat{p}_{a_{2}}$ and $r_{a_{3}}=\hat{p}_{a_{3}}$. Trivially, $a_{1}$ has no incentive to change $d_{a_{1}}$. On the other hand, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$. This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\hat{p}_{a_{2}}$ and $\hat{p}_{a_{3}}$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$. Then, in this case, $\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right)$ is a Nash equilibrium. Thus, a wasteful assignment can occur as a Nash equilibrium of agents’ strategies when the assignment is determined by the uniform rank-minimizing mechanism.
@@
Suppose that the uniform rank-minimizing mechanism used. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
@@
This result implies that if there is $a\in\mathcal{A}$ satisfying (3) for some $o,o^{\prime}\in\mathcal{O}$, then agent $a$ has a strong incentive to adopt an ODS. This is satisfied only when some agent has at least one unacceptable object. The condition given in (3) is that there is a truly unacceptable type $o^{\prime}$ such that the number of copies of $o^{\prime}$, plus the sum of the copies of the truly acceptable types, is sufficiently small. We discuss the cases where (3) is satisfied for some $a,o,o^{\prime}$ in the subsequent section. Note that there is $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $R\left(\varnothing,p_{a^{\prime}}\right)=\left|\mathcal{O}\right|$ for all $a^{\prime}\in\mathcal{A\setminus}\left\{a\right\}$ and $a$ strictly prefers $g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)$ to $g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)$. Thus, for all $a^{\prime}$, $p_{a^{\prime}}$ may also be an ODS of $a^{\prime}$.
@@
As stated earlier, there may exist multiple ODSs. In that case, an ODS may be strategically dominated by another one. We have the following result.
Proposition 3
$R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)<R\left(o^{\prime},r_{a}\right)\text{ and }\sum\limits_{\bar{o}:R\left(\bar{o},r_{a}\right)<R\left(\varnothing,r_{a}\right)}q_{\bar{o}}+q_{o^{\prime}}\geq\left|\mathcal{A}\right|,$
for some $o,o^{\prime}\in\mathcal{O}$. Let $d_{a}\in\mathcal{D}_{a}$ be such that $R\left(\varnothing,r_{a}\right)=R\left(o^{\prime},d_{a}\right)$. Then, for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and all $\bar{o}\in\mathcal{O}$,
$g\left(f^{U}\left(d_{a},p_{-a}\right),r\right)_{a\bar{o}}=g\left(f^{U}\left(r_{a},p_{-a}\right),r\right)_{a\bar{o}}.$
In this case, truth-telling is not strategically dominated by an ODS of revealing such $d_{a}$. Note that revealing the full extension may be such an ODS, when $o^{\prime}$ defined in Proposition 3 satisfies $R\left(\varnothing,r_{a}\right)+1=R\left(o^{\prime},r_{a}\right)$.
Moreover, by Theorem 1 and Proposition 3, we show that for an agent, revealing the full extension may be strategically dominated by another ODS. First, we provide an example to briefly see this fact.

### v5

@@ 3 Rank-minimizing Mechanism
otherwise; that is, if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\alpha^{\prime},\alpha,\alpha\right)$ nor $\left(\alpha,\alpha^{\prime},\alpha\right)$ nor $\left(\alpha,\alpha,\alpha^{\prime}\right)$. In words, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00.
We additionally assume that if $\left(p_{a_{i}},p_{a_{j}},p_{a_{k}}\right)=\left(\alpha^{\prime},\alpha,\alpha\right)$, then
$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{j}\bar{o}}=\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{k}\bar{o}}$
for all $\bar{o}\in\mathcal{O}$. Then, $f^{\ast}$ satisfies weak ETE. However, whether $f^{\ast}$ satisfies ETE or not is dependent on $q_{o_{2}}$.
@@
We assume that $r_{a_{1}}=\alpha^{\prime}$ and this rank-minimizing mechanism $f^{\ast}$ is employed. First, suppose $q_{o_{2}}=1$. In this case, $f^{\ast}$ satisfies ETE and thus this is a fair rank-minimizing mechanism. In this case, $\bar{k}\left(\alpha\right)=\bar{k}\left(\alpha^{\prime}\right)=3$. Then, for example, $r_{a_{1}}=\alpha^{\prime}$ is not strategically dominated by $\alpha$ and any other preference orders, because
$\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f^{\ast}\left(\alpha^{\prime},\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{3}}=\frac{1}{3},$
$\left(f^{\ast}\left(\alpha,\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{1}}$ $=$ $\left(f^{\ast}\left(\alpha,\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}o_{2}}=\left(f^{\ast}\left(\alpha,\alpha^{\prime},\alpha^{\prime}\right)\right)_{a_{1}\varnothing}=\frac{1}{3}.$
That is, if $p_{a_{2}}=p_{a_{3}}=\alpha^{\prime}$, then $a_{1}$ is worse off by revealing $\alpha$ instead of $\alpha^{\prime}=r_{a_{1}}\,$(truth-telling).
@@
Second, suppose $q_{o_{2}}\geq 2$. In this case, $\bar{k}\left(\alpha\right)=\bar{k}\left(\alpha^{\prime}\right)=2$. Moreover, if $p_{a_{i}}$ is either $\alpha$ or $\alpha^{\prime}$, then a rank-minimizing mechanism satisfies
$\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{i}o_{1}}+\left(f^{\ast}\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)\right)_{a_{i}o_{2}}=1.$
@@
That is, if a rank-minimizing mechanism is employed and $a_{1}$ chooses either $\alpha$ or $\alpha^{\prime}$, then $a_{1}$ must have either $o_{1}$ or $o_{2}$. Thus, the only strategic interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. If $f^{\ast}$ is employed, then $\alpha^{\prime}=r_{a_{1}}$ is strategically dominated by $\alpha$ for $a_{1}$ due to (2). Therefore, we conclude that even if a rank-minimizing mechanism that satisfies weak ETE is employed, truth-telling may be strategically dominated by some other strategy. Note that in the same case, $f^{\ast}$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\alpha$ and $\alpha^{\prime}$ are essentially the same in this case. However, $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{2}}>0$ or $\left(f^{\ast}\left(\alpha^{\prime},\alpha,\alpha\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Hence $f^{\ast}$ does not satisfy ETE.
@@ 4 Refusal option
This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost other previous studies related to ours. [^4] As exceptions, Feigenbaum et al. (2020) and Afacan (2022) consider an assignment problem where, after the initial allocation is decided, if someone refuses the assigned object, it is reallocated. Moreover, Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
@@
We reconsider strategic problems of a fair rank-minimizing mechanism in this three-stage game. For a mechanism $f$ and an agent $a$, (revealing) $p_{a}^{\prime}$ is said to be weakly strategically dominated by (revealing) $p_{a}$ with the refusal option if $a$ weakly prefers $g\left(f\left(p_{a},p_{-a}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}\right),r\right)$ for all $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$. Moreover, for $f$ and $a$, (revealing) $p_{a}^{\prime}$ is said to be (strictly) strategically dominated by (revealing) $p_{a}$ with the refusal option if $p_{a}^{\prime}$ is weakly strategically dominated by $p_{a}$ with the refusal option and there exists some $p_{-a}^{\prime}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ such that $a$ strictly prefers $g\left(f\left(p_{a},p_{-a}^{\prime}\right),r\right)$ to $g\left(f\left(p_{a}^{\prime},p_{-a}^{\prime}\right),r\right)$.
We show that, under the uniform rank-minimizing mechanism, truth-telling is strategically dominated by another strategy when the agents have the refusal option. We define the following specific strategies. Let $\mathcal{D}_{a}\subseteq\mathcal{P}$ be a set of preferences such that $d_{a}\in\mathcal{D}_{a}$ if (i) $R\left(o,r_{a}\right)=R\left(o,d_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$ and (ii) $R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$. First, the ranks of the truly acceptable object types are unchanged. Second, the null object type $\varnothing$ is the least preferable one under $d_{a}\in\mathcal{D}_{a}$. We call the strategy of agent $a$ of revealing a preference in $\mathcal{D}_{a}$ an outside option demotion strategy (hereafter ODS) of agent $a$. Under an ODS, every object type that is ranked below $\varnothing$ in the agent’s true preference is instead ranked above $\varnothing$ in $d_{a}$. Note that an ODS (revealing $d_{a}\in\mathcal{D}_{a}$) increases the number of acceptable object types revealed by $a$; that is, $R\left(\varnothing,r_{a}\right)\leq R\left(\varnothing,d_{a}\right)=\left|\mathcal{O}\right|$ is satisfied.
@@
Let $\mathcal{A=}\left\{a_{1},a_{2},a_{3}\right\}$ and $\mathcal{O=}\left\{o_{1},o_{2},\varnothing\right\}$. Let $\beta=\left(o_{1},o_{2},\varnothing\right)$, $\beta^{\prime}=\left(o_{2},o_{1},\varnothing\right)$ and $\beta^{\prime\prime}=(o_{1},\varnothing,o_{2})$. Suppose $q_{o_{1}}=q_{o_{2}}=1$ and $r_{a_{1}}=\beta^{\prime\prime}$. Then, the unique ODS is revealing $d_{a_{1}}=\beta$. In this example, we assume $\left(p_{a_{2}},p_{a_{3}}\right)=\left(r_{a_{2}},r_{a_{3}}\right);$ that is, $a_{2}$ and $a_{3}$ always adopt truth-telling.
First, we consider the case where $p_{a_{2}}=\beta$ and $p_{a_{3}}=\beta^{\prime}$. Then,
$f^{U}\left(r_{a_{1}},\beta,\beta^{\prime}\right)=y=\left(\begin{array}[]{ccc}y_{a_{1}o_{1}}&y_{a_{1}o_{2}}&y_{a_{1}\varnothing}\\ y_{a_{2}o_{1}}&y_{a_{2}o_{2}}&y_{a_{2}\varnothing}\\ y_{a_{3}o_{1}}&y_{a_{3}o_{2}}&y_{a_{3}\varnothing}\end{array}\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right).$
@@
$f^{U}\left(d_{a_{1}},\beta,\beta^{\prime}\right)=\left(\begin{array}[]{ccc}1/2&0&1/2\\ 1/2&0&1/2\\ 0&1&0\end{array}\right)$
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$. First, under $\left(r_{a_{1}},\beta,\beta^{\prime}\right),$varnothing is ranked lowest by $a_{2}$ and $a_{3}$ but $a_{1}$ ranks $\varnothing$ as their second-best. Thus, in any rank-minimizing mechanism, $a_{1}$ is assigned $\varnothing$ with probability 1. On the other hand, since $a_{1}$ places $\varnothing$ at the bottom of their ranking. As a result. under $\left(d_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$, $a_{1}$ is assigned $o_{1}$—the object that is truly their best—with probability $1/2$, and $\varnothing$ with probability $1/2$. Thus, in this case, $a_{1}$ is better off by choosing $d_{a_{1}}$ instead of $r_{a_{1}}$.
Second, suppose $p_{a_{2}}=p_{a_{3}}=\beta^{\prime\prime}.$ Then,
$f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right),$
$f^{U}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)=\left(\begin{array}[]{ccc}1/3&2/3&0\\ 1/3&0&2/3\\ 1/3&0&2/3\end{array}\right).$
Thus, in this case, $a_{1}$ seems to be worse off by revealing $d_{a_{1}}$ instead of $r_{a_{1}}$, because $a_{1}$ prefers $\varnothing$ to $o_{2}$. However, $a_{1}$ can also refuse $o_{2}$ and obtain $\varnothing$. Therefore, if $r_{a_{2}}=r_{a_{3}}=\beta^{\prime\prime}$,
$g\left(f^{U}\left(d_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right),r\right)=g\left(f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right),r\right)=f^{U}\left(r_{a_{1}},\beta^{\prime\prime},\beta^{\prime\prime}\right)\text{;}$
@@
Thus, in this example, $a_{1}$ is never worse off and may even be better off by selecting $d_{a_{1}}$ (the ODS) instead of $r_{a_{1}}$. Next, we show that the ODS by $a_{1}$ may make the assignment wasteful.
Third, suppose $p_{a_{2}}=p_{a_{3}}=\beta$. Then,
$f^{U}\left(r_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}0&0&1\\ 1/2&1/2&0\\ 1/2&1/2&0\end{array}\right),$
$f^{U}\left(d_{a_{1}},\beta,\beta\right)=\left(\begin{array}[]{ccc}1/3&1/3&1/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
In the latter case, $a_{1}$ refuses the assignment; that is,
$g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/3&1/3\\ 1/3&1/3&1/3\end{array}\right).$
Then, $a_{1}$ is better off by adopting the ODS instead of truth-telling. However, $g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)$ is wasteful, because
$\sum\limits_{a\in\left\{a_{1},a_{2},a_{3}\right\}}g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{ao_{2}}$ $=$ $\frac{2}{3}<1,$
and $\text{and }g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)_{a^{\prime}\varnothing}$ $=$ $\frac{1}{3}>0,$
for $a^{\prime}=a_{2},a_{3}$. That is, for
@@
$a_{2}$ and $a_{3}$ strictly prefer and $a_{1}$ weakly prefers $\tilde{x}$ to $g\left(f^{U}\left(d_{a_{1}},\beta,\beta\right),r\right)$.
If $a_{1}$ is assigned to $o_{2}$, then either $a_{2}$ or $a_{3}$ will be assigned to $\varnothing$. However, in this case, $a_{1}$ will refuse $o_{2}$ and thus $o_{2}$ will be remained unallocated. In this case, rather than leaving $o_{2}$ unallocated, it is more efficient to assign $o_{2}$ to the agent ($a_{2}$ or $a_{3}$) who would otherwise receive $\varnothing$ and instead assign $a_{1}$ to $\varnothing$ from the beginning. This adjustment ensures that $o_{2}$ is allocated efficiently. Hence revealing $d_{a_{1}}$ and refusing an unacceptable assignment may benefit $a_{1}$ but lead to a wasteful allocation, as some agents may prefer the object type that $a_{1}$ refuses.
Note that $\left(d_{a_{1}}\left(=\beta\right),\beta,\beta\right)$ can be a Nash equilibrium. Trivially, $a_{1}$ has no incentive to change $d_{a_{1}}$. On the other hand, suppose that both $a_{2}$ and $a_{3}$ prefer the assignment such that each object $o\in\left\{o_{1},o_{2},\varnothing\right\}$ assigns with the probability $1/3$ to the assignment such that $o_{2}$ assigns with probability $1$. This ensures that neither $a_{2}$ nor $a_{3}$ has an incentive to change $\beta$ and $\beta$ to $(o_{2},o_{1},\varnothing)$ or $(o_{2},\varnothing,o_{1})$. Then, in this case, $\left(d_{a_{1}},\beta,\beta\right)$ is a Nash equilibrium. Thus, a wasteful assignment can occur as a Nash equilibrium of agents’ strategies when the assignment is determined by the uniform rank-minimizing mechanism.
@@
Suppose that the uniform rank-minimizing mechanism is employed. Then, for all agents, truth-telling is weakly strategically dominated by any ODS with the refusal option.
@@
This result implies that if there is $a\in\mathcal{A}$ satisfying (3) for some $o,o^{\prime}\in\mathcal{O}$, then agent $a$ has a strong incentive to adopt an ODS. This is satisfied only when some agent has at least one unacceptable object. The condition given in (3) is that there is a truly unacceptable type $o^{\prime}$ such that the number of copies of $o^{\prime}$, plus the sum of the copies of the truly acceptable types, is sufficiently small. We discuss real-world scenarios where (3) is satisfied for some $a,o,o^{\prime}$ in the subsequent section.
@@
As stated earlier, multiple ODSs may exist. Even if an ODS strategically dominates truth-telling, another ODS that does not strategically dominate truth-telling may also exist. In such a case, the latter ODS is strategically dominated by the former. The strategy of revealing the full extension is not an exception to this. We show this fact by introducing the following example.